Problem 12-5A Partner withdrawal and admission LO P3, P4

[The following information applies to the questions displayed below.]

Meir, Benson, and Lau are partners and share income and loss in a 3:2:5 ratio (in percents: Meir, 30%; Benson, 20%; and Lau, 50%). The partnership's capital balances are as follows: Meir, $88,000; Benson, $59,000; and Lau, $153,000. Benson decides to withdraw from the partnership.

Problem 12-5A Part 1

1. Prepare the journal entry to record Benson's withdrawal under each independent assumptions. (Do not round intermediate calculations.)

(a) Benson sells her interest to North for $160,000 after North is approved as a partner; (b) Benson gives her interest to a son-in-law, Schmidt, and Schmidt is approved as a partner; (c) Benson is paid $59,000 in partnership cash for her equity; (d) Benson is paid $97,000 in partnership cash for her equity; and (e) Benson is paid $15,000 in partnership cash plus equipment recorded on the partnership books at $35,000 less its accumulated depreciation of $11,600.

Record the admission of Rhode with an investment of $100,000 for a 25% interest in the equity.

Record the admission of Rhode with an investment of $73,000 for a 25% interest in the equity.

Record the admission of Rhode with an investment of $131,000 for a 25% interest in the equity.
Answer #1
Event General journal Debit Credit
A) Benson's capital $ 59,000
                  North's capital $ 59,000
( To record Benson sells his interest to North )
B) Benson's capital $ 59,000
               Schmidt's Capital $ 59,000
( To record Benson gives his interest to Schmidt )
C) Benson's capital $ 59,000
                               Cash     $ 59,000
(To record Benson's paid in partnership for the equity)
D) Benson's capital $ 59,000
Meir's capital   (see note 1 ) $ 14,250
Lau's capital     (see note 1 ) $ 23,750
                                    Cash $ 97,000
(To record Benson's paid in partnership for the equity in the ratio 3:5)
E) Benson's capital $ 59,000
Accumulated Depreciation-Equipment $ 11,600
                       Equipment $ 35,000
                       Cash $ 15,000
                      Meir's capital   (see note 2 ) $ 7,725
                      Lau's capital     (see note 2 ) $ 12,875
(To record Benson's paid in partnership plus equipment
less accumulated depreciation will be share in 3 : 5 ratio )
REQUIRED 2 :
Event General journal Debit Credit
1) Cash $ 100,000
                 Rhode's Capital (see note 3 ) $ 100,000
(to record admission of Rhode with an investment of $ 100,000 for 25% interest in equity)
2) Cash $ 73,000
Meir's capital   (see note 4) $ 6,075
Benson's capital ( see note 4 ) $ 4,050
Lau's capital     (see note 4 ) $ 10,125
                          Rhode's Capital (see note 4 ) $ 93,250
(to record admission of Rhode with an investment of $ 73,000 for 25% interest in equity)
3) Cash $ 131,000
                                   Meir's capital   (see note 5) $ 6,975
                                     Benson's capital ( see note 5 ) $ 4,650
                                       Lau's capital     (see note 5 ) $ 11,625
                                       Rhode's Capital (see note 5 ) $ 107,750
(to record admission of Rhode with an investment of $ 131,000 for 25% interest in equity)

EXPLANATION :

Note1 :
Meir's capital = ($97,000 - $59,000)* 3/ 8 = $ 14,250
Lau's capital = ($97,000 - $59,000)* 5/ 8 = $ 23,750
So, Remaining capital after Benson will be shared
in the ratio 3 : 5
Note 2 :
Meir's capital = {($ 59,000 - ($ 35,000 -$ 11,600 +$ 15,000)}* 3 /8 = $ 7,725
Lau's capital ={($ 59,000 - ($ 35,000 -$ 11,600 +$ 15,000)}* 5/8 = $ 12,875
So, Remaining capital after Benson will be shared
in the ratio 3 : 5
Note 3 :
To calculate Rhodes capital :
Total capital ($ 88,000 + $ 59,000 + $ 153,000 )= $ 300,0000
(Total capital + investment )*25 %
($300,000 + $ 100,000 )*25 % = $ 100,000
Thus , No bonus is to be received or paid
Note 4 :
Total capital ($ 88,000 + $ 59,000 + $ 153,000 )= $ 300,0000
(Total capital + investment )*25 %
($300,000 + $ 73,000 )*25 % = $ 93,250
($ 73,000 - $ 93,250 ) = $ (20,250 )
Above value to be shared in the ratio of 3:2:5 for remaining partner
Meir's capital = 20,250 *3/10 = $ 6,075
Benson's Capital = 20,250*2/10= $ 4,050
Lau's Capital = 20,250 * 5/10 = $ 10,125
Hence, bonus paid to be new partner ( Rhode)
Note 5 :
Total capital ($ 88,000 + $ 59,000 + $ 153,000 )= $ 300,0000
(Total capital + investment )*25 %
($300,000 + $ 131,000 )*25 % = $ 107,750
($ 131,000 - $ 107,750 ) = $ 23,250
Above value to be shared in the ratio of 3:2:5 for remaining partner
Meir's capital = 23,250 *3/10 = $ 6,975
Benson's Capital = 23,250*2/10= $ 4,650
Lau's Capital = 23,250 * 5/10 = $ 11,625
Hence, bonus is to be received by Old partners